# Physically Valid Biomolecular Interaction Modeling with Gauss-Seidel Projection

### v1

@@ Appendix B Additional Qualitative Results
Appendix C Training and Inference Algorithms
We provide pseudo-code for the training and inference algorithms of our model, shown in Algorithms 3 and 4. The values of hyperparameters shown in both algorithms follow (Gong et al., 2025).
@@ Appendix C Training and Inference Algorithms
Appendix D Proof for the Convergence of Linearization
Theorem D.1
@@ Appendix D Proof for the Convergence of Linearization
with $\mathbf{K}=\mathbf{I}-\sum_{j}\lambda_{j}\nabla^{2}\mathbf{C}_{j}(\mathbf{x}^{(i)})$. We thus solve, at each iteration, a linear system obtained by Newton linearization of the coupled equations. By Newton’s local convergence theorem (Sauer, 2011), the iterates converge quadratically given the Jacobian at the solution $(\mathbf{x}^{\ast},\bm{\lambda}^{\ast})$ is non-singular and the initialization lies in its basin of attraction. Taking the Schur complement (Zhang, 2005) with respect to $\bm{\alpha}\!\succ\!\mathbf{0}$ shows that the non-singularity of the Jacobian is equivalent to the invertibility of
@@
Appendix E Proof for the Convergence and Constraint Satisfaction of Gauss-Seidel Projection
Theorem E.1
@@ Appendix E Proof for the Convergence and Constraint Satisfaction of Gauss-Seidel Projection
A cyclic Gauss-Seidel sweep on Eq. 10 is exactly cyclic coordinate descent on Eq. 12: each update minimizes $f$ w.r.t. a single coordinate $\Delta\lambda_{j}$ while holding the others fixed, yielding the closed form in Eq. 11. Because $f$ is strongly convex and quadratic, each coordinate step strictly decreases $f$ unless already optimal, and the sequence $\{f(\Delta\bm{\lambda}^{(k)})\}$ is monotonically decreasing. For SPD systems, both classical matrix-splitting analysis and coordinate-descent theory guarantee linear convergence of cyclic Gauss-Seidel to the unique minimizer (Varga, 1962; Saad, 2003; Wright, 2015). In particular, there exists $\rho\in(0,1)$ such that $\|\Delta\bm{\lambda}^{(k)}-\Delta\bm{\lambda}^{\ast}\|_{\mathbf{A}}\leq\rho^{k}\|\Delta\bm{\lambda}^{(0)}-\Delta\bm{\lambda}^{\ast}\|_{\mathbf{A}}$. Each coordinate update accesses only the atoms affected by the current constraint, so it is $O(1)$. A full sweep is thus $O(m)$.
The proof of Theorem D.1 shows that Eq. 10 is a quasi-Newton approximation of the first-order optimality condition Eq. 3. Hence, the convergence speed of the Gauss-Seidel projection is of quasi-second order. With a strict physical tolerance and sufficiently small penalty parameter $\bm{\alpha}$, the final iterate satisfies the optimality conditions of the penalized projection, meaning that in practice it realizes a hard-constrained projection (all constraints are strictly satisfied) (Boyd & Vandenberghe, 2004).
Appendix F Proof for the Convergence of Conjugate Gradient
Theorem F.1
@@ Appendix F Proof for the Convergence of Conjugate Gradient
This matrix is SPD, and so is $\mathbf{H}(\mathbf{x}_{\mathrm{proj}})+\mathbf{I}$. The conjugate gradient method is guaranteed to converge for any linear system with a SPD coefficient matrix (Sauer, 2011), and its convergence rate is governed by the condition number of the matrix. Since the coefficient matrix contains the identity $\mathbf{I}$, the condition number is bounded below by the size of $\mathbf{I}$, i.e., $m$, which ensures that the convergence of conjugate gradient remains stable in the neighborhood of feasibility.
Appendix G Detailed Formulation of Physical Validity Constraints
@@
Below we present the formalization of each of the physical constraints following Boltz-1-Steering (Wohlwend et al., 2024), which are visualized in Fig. 8.
@@
where $L$ and $U\in\mathbb{R}^{N\times N}$ denote the lower and upper bounds matrices. For a pair of atoms $(i,j)$, $L_{i,j}$ gives the lower distance bound and $U_{i,j}$ gives the upper distance bound (Buttenschoen et al., 2024).
@@
Appendix H Details on Baselines and Evaluation Protocol
We compare our method against strong baselines: Boltz-1 (the original model) (Wohlwend et al., 2024), Boltz-1-Steering (Boltz-1 with FK-steering at inference), Boltz-2 (the updated Boltz model) (Passaro et al., 2025), Protenix (Team et al., 2025) and Protenix-Mini (Gong et al., 2025). All baselines use 200-step diffusion sampling, whereas Protenix-Mini uses a 5-step sampling. We evaluate our method and all baselines on six benchmarks: CASP15, Test, PoseBusters, AF3-AB, dsDNA, and RNA-Protein. CASP15 and Test (from Boltz-1 (Wohlwend et al., 2024)) cover diverse assemblies, including protein complexes, nucleic acids, and small molecules. PoseBusters (Buttenschoen et al., 2024) focuses on protein-ligand modeling. AF3-AB (released with AlphaFold3 (Abramson et al., 2024)) contains protein-antibody complexes. dsDNA (Ma et al., 2025) contains entries with two DNA chains and one protein chain. RNA-Protein (Ma et al., 2025) contains entries consisting of one RNA chain and one protein chain.
We compare our method with baselines using standard evaluation metrics, including Complex LDDT, Prot-Prot LDDT, Lig-Prot LDDT, DNA-Prot LDDT, RNA-Prot LDDT, DockQ > 0.23, Mean LDDT+PLI, L-RMSD < 2 Å, and Physical Validity. Complex LDDT measures the overall geometric accuracy throughout the complex, while Prot-Prot LDDT, Lig-Prot LDDT, DNA-Prot LDDT, and RNA-Prot LDDT focus specifically on the quality of protein-protein, protein-ligand, protein-DNA, and protein-RNA interfaces, respectively. The average DockQ success rate (DockQ > 0.23), defined as the proportion of predictions with DockQ > 0.23, measures the fraction of cases where good protein-protein interactions are correctly predicted. The average protein-ligand interface LDDT (Mean LDDT+PLI) combines structural accuracy (LDDT) with protein-ligand interaction scores (PLI) to provide a balanced assessment of both geometry and intermolecular contacts. L-RMSD is defined as the proportion of ligands with a pocket-aligned RMSD below $2$ Å, a widely adopted measure of molecular docking accuracy. Finally, Physical Validity checks whether the predicted structures obey fundamental physical constraints, including steric clash, tetrahedral atomic chirality, bond stereochemistry, planar double bonds, internal geometry, overlay chains, and covalently bound chains. All evaluation metrics follow those used in Boltz-1 (Wohlwend et al., 2024) and Protenix (Team et al., 2025).
Appendix I The Use of Large Language Models (LLMs)

### v2

@@ Appendix G Analysis of Structural Perturbations Introduced by Gauss-Seide Projection
Appendix H Training and Inference Algorithms
We provide pseudo-code for the training and inference algorithms of our model, shown in Algorithms 3 and 4. The values of hyperparameters shown in both algorithms follow (gong2025protenixminiefficientstructurepredictor).
@@ Appendix H Training and Inference Algorithms
Appendix I Proof for the Convergence of Linearization
Theorem I.1
@@ Appendix I Proof for the Convergence of Linearization
with $\mathbf{K}=\mathbf{I}-\sum_{j}\lambda_{j}\nabla^{2}\mathbf{C}_{j}(\mathbf{x}^{(i)})$. We thus solve, at each iteration, a linear system obtained by Newton linearization of the coupled equations. By Newton’s local convergence theorem (sauer2011numerical), the iterates converge quadratically given the Jacobian at the solution $(\mathbf{x}^{\ast},\bm{\lambda}^{\ast})$ is non-singular and the initialization lies in its basin of attraction. Taking the Schur complement (fuzhen2005schur) with respect to $\bm{\alpha}\!\succ\!\mathbf{0}$ shows that the non-singularity of the Jacobian is equivalent to the invertibility of
@@
Appendix J Proof for the Convergence and Constraint Satisfaction of Gauss-Seidel Projection
Theorem J.1
@@
A cyclic Gauss-Seidel sweep on Eq. 10 is exactly cyclic coordinate descent on Eq. 12: each update minimizes $f$ w.r.t. a single coordinate $\Delta\lambda_{j}$ while holding the others fixed, yielding the closed form in Eq. 11. Because $f$ is strongly convex and quadratic, each coordinate step strictly decreases $f$ unless already optimal, and the sequence $\{f(\Delta\bm{\lambda}^{(k)})\}$ is monotonically decreasing. For SPD systems, both classical matrix-splitting analysis and coordinate-descent theory guarantee linear convergence of cyclic Gauss-Seidel to the unique minimizer (varga1962iterative; saad2003iterative; wright2015coordinate). In particular, there exists $\rho\in(0,1)$ such that $\|\Delta\bm{\lambda}^{(k)}-\Delta\bm{\lambda}^{\ast}\|_{\mathbf{A}}\leq\rho^{k}\|\Delta\bm{\lambda}^{(0)}-\Delta\bm{\lambda}^{\ast}\|_{\mathbf{A}}$. Each coordinate update accesses only the atoms affected by the current constraint, so it is $O(1)$. A full sweep is thus $O(m)$.
The proof of Theorem I.1 shows that Eq. 10 is a quasi-Newton approximation of the first-order optimality condition Eq. 3. Hence, the convergence speed of the Gauss-Seidel projection is of quasi-second order. With a strict physical tolerance and sufficiently small penalty parameter $\bm{\alpha}$, the final iterate satisfies the optimality conditions of the penalized projection, meaning that in practice it realizes a hard-constrained projection (all constraints are strictly satisfied) (boyd2004convex).
Appendix K Proof for the Convergence of Conjugate Gradient
Theorem K.1
@@
This matrix is SPD, and so is $\mathbf{H}(\mathbf{x}_{\mathrm{proj}})+\mathbf{I}$. The conjugate gradient method is guaranteed to converge for any linear system with a SPD coefficient matrix (sauer2011numerical), and its convergence rate is governed by the condition number of the matrix. Since the coefficient matrix contains the identity $\mathbf{I}$, the condition number is bounded below by the size of $\mathbf{I}$, i.e., $m$, which ensures that the convergence of conjugate gradient remains stable in the neighborhood of feasibility.
Appendix L Detailed Formulation of Physical Validity Constraints
@@
Below we present the formalization of each of the physical constraints following Boltz-1-Steering (wohlwend2024boltz1), which are visualized in Fig. 11.
@@
where $L$ and $U\in\mathbb{R}^{N\times N}$ denote the lower and upper bounds matrices. For a pair of atoms $(i,j)$, $L_{i,j}$ gives the lower distance bound and $U_{i,j}$ gives the upper distance bound (buttenschoen2024posebusters).
@@
Appendix M Details on Baselines and Evaluation Protocol
We compare our method against strong baselines: Boltz-1 (the original model) (wohlwend2024boltz1), Boltz-1-Steering (Boltz-1 with FK-steering at inference), Boltz-2 (the updated Boltz model) (passaro2025boltz2), Protenix (bytedance2025protenix) and Protenix-Mini (gong2025protenixminiefficientstructurepredictor). All baselines use 200-step diffusion sampling, whereas Protenix-Mini uses a 5-step sampling. We evaluate our method and all baselines on six benchmarks: CASP15, Test, PoseBusters, AF3-AB, dsDNA, and RNA-Protein. CASP15 and Test (from Boltz-1 (wohlwend2024boltz1)) cover diverse assemblies, including protein complexes, nucleic acids, and small molecules. PoseBusters (buttenschoen2024posebusters) focuses on protein-ligand modeling. AF3-AB (released with AlphaFold3 (AlphaFold3)) contains protein-antibody complexes. dsDNA (PXMeter) contains entries with two DNA chains and one protein chain. RNA-Protein (PXMeter) contains entries consisting of one RNA chain and one protein chain.
We compare our method with baselines using standard evaluation metrics, including Complex LDDT, Prot-Prot iLDDT, Lig-Prot iLDDT, DNA-Prot iLDDT, RNA-Prot iLDDT, DockQ > 0.23, L-RMSD < 2 Å, TM-score, and Physical Validity. Complex LDDT measures the overall geometric accuracy throughout the complex, while Prot-Prot iLDDT, Lig-Prot iLDDT, DNA-Prot iLDDT, and RNA-Prot iLDDT focus specifically on the quality of protein-protein, protein-ligand, protein-DNA, and protein-RNA interfaces, respectively. The average DockQ success rate (DockQ > 0.23), defined as the proportion of predictions with DockQ > 0.23, measures the fraction of cases where good protein-protein interactions are correctly predicted. L-RMSD is defined as the proportion of ligands with a pocket-aligned RMSD below $2$ Å, a widely adopted measure of molecular docking accuracy. TM-score is a length-normalized measure of structural similarity, widely used to assess protein structure prediction accuracy. Finally, Physical Validity checks whether the predicted structures obey fundamental physical constraints, including steric clash, tetrahedral atomic chirality, bond stereochemistry, planar double bonds, internal geometry, overlapping chains, and covalently bound chains. All evaluation metrics follow those used in Boltz-1 (wohlwend2024boltz1) and Protenix (bytedance2025protenix).
Appendix N The Use of Large Language Models (LLMs)